OpenStudy (anonymous):

Nernoulli equation y'+6xy = 3xy^3, y(0)=5 what is the answer for this one??

OpenStudy (inkyvoyd):

Use wolfram alpha if you want to find the answer. Use wolfram alpha show steps if you want the process. LOL.

OpenStudy (anonymous):

Make a substitution of u=y^(1-3) which makes it linear See http://www.math.oregonstate.edu/home/programs/undergrad/CalculusQuestStudyGuides/ode/first/bernoulli/bernoulli.html

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

\[y'+6xy=3xy^3\] \[v=y^{1-3}=y^{-2}\\ v'=-2y^{-3}y'\\ y'=-\frac{1}{2}y^3v'\] \[-\frac{1}{2}y^3v'+6xy=3xy^3\\ -\frac{1}{2}v'+6xy^{-2}=3x\\ -\frac{1}{2}v'+6xv=3x\\ v'-12xv=-6x\] Now you've got yourself a linear ODE. Find the integrating factor and solve.